To prevent tank rupture during deep-space travel, an engineering team is studying the effect of temperature on gases confined to small volumes. What is the pressure of 2.60 mol of gas D measured at 251 ∘C in a 1.75-L container assuming real behavior?
Answer #1

For real gas, the modified ideal gas law is,

(P + an²/V²) x (V - nb) = nRT

Therefore,

P = nRT/(V - nb) - an²/V²

Where,

V = 1.75 L

R = 0.08206 L atm/K mol

T = 251°C = 524.15 K

n = 2.40 moles

a = 4.16 L² atm/mol²

b = 0.0305 L/mol

By substituting the values,

P = (2.4 x 0.08206 x 524.15 )/(1.75 -2.4 x 0.0305) – (4.16 x 2.42) / 1.752

P = (103.228/1.6768) –(23.96/3.0625)

= 61.5625 – 7.8236 = 53.7389 atm

P = 53.7389 atm